OpenStudy (anonymous):

Simplify. 3 square root of 5 end root minus 2 square root of 2 end root plus square root of 45 end root minus square root of 28 2 square root of 12 2 square root of 2 6 square root of 5 end root minus 4 square root of 7 6 square root of 10 end root minus 4 square root of 14

OpenStudy (anonymous):

@Mokeira

OpenStudy (mokeira):

\[3\sqrt{5}-2\sqrt{2}+\sqrt{45}\] first try and simplify\[\sqrt{45}\] tell me what you get

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

Can you explain what I have to do to simplify?

OpenStudy (mokeira):

try to think of two number you would multiply to get 45

OpenStudy (anonymous):

9 times 5

OpenStudy (mokeira):

exactly so \[\sqrt{9\times5}\] simplify